# Learning to Diagnose and Correct Moral Errors: Towards Enhancing Moral Sensitivity in Large Language Models

### v1

@@ 4.2 Experimental Results
In this section, we present our experimental results across different benchmarks, which contain different types of moral discourses that vary in linguistic expressions. This allows us to analyze how the light- and heavy-load pragmatic inference methods perform with each benchmark. Additionally, we report the experimental results that the diagnostic outputs from learning2DiagCorr can significantly improve the moral quality of responses generated by off-the-shelf LLMs through direct prompting. This pertains to extrinsic self-correction liu2024self, which is particularly challenging for LLMs that have not been exposed to the training data. Thus, strong performance in correcting moral errors reflects the high quality and effectiveness of the diagnoses produced by learning2DiagCorr.
@@ A.1 Dataset Description
Direct Toxic Language: To prepare the training dataset, we sample 2,000 prompts $x_{p}$ and their corresponding continuations $x_{r}$ from the RealToxicityPrompts benchmark gehman2020realtoxicityprompts.In this dataset, half of the replies have a toxicity score below 0.1, and the other half have a toxicity score above 0.8, where toxicity is evaluated using the PerspectiveAPI. For continuations $x_{r}$ with a toxicity score greater than 0.8, we use DeepSeek to revise $x_{r}$ into a corrected version $y_{r}$ whose toxicity score is below 0.1. All toxicity scores are evaluated using the PerspectiveAPI. To construct the testing dataset, We sample the data from the testing dataset in RealToxicityPrompts where the prompts has the toxic level < 0.1 and its continuation has toxicity score greater than 0.8.

### v2

@@ 4.2 Experimental Results
In this section, we present our experimental results across different benchmarks, which contain different types of moral discourses that vary in linguistic expressions. This allows us to analyze how the light- and heavy-load pragmatic inference methods perform with each benchmark. Additionally, we report the experimental results that the diagnostic outputs from learning2DiagCorr can significantly improve the moral quality of responses generated by off-the-shelf LLMs through direct prompting. This pertains to extrinsic self-correction Liu et al. (2024b), which is particularly challenging for LLMs that have not been exposed to the training data. Thus, strong performance in correcting moral errors reflects the high quality and effectiveness of the diagnoses produced by learning2DiagCorr.
@@ A.1 Dataset Description
Direct Toxic Language: To prepare the training dataset, we sample 2,000 prompts $x_{p}$ and their corresponding continuations $x_{r}$ from the RealToxicityPrompts benchmark Gehman et al. (2020).In this dataset, half of the replies have a toxicity score below 0.1, and the other half have a toxicity score above 0.8, where toxicity is evaluated using the PerspectiveAPI. For continuations $x_{r}$ with a toxicity score greater than 0.8, we use DeepSeek to revise $x_{r}$ into a corrected version $y_{r}$ whose toxicity score is below 0.1. All toxicity scores are evaluated using the PerspectiveAPI. To construct the testing dataset, We sample the data from the testing dataset in RealToxicityPrompts where the prompts has the toxic level < 0.1 and its continuation has toxicity score greater than 0.8.